What is the exact value of the expression the square root of 486 − the square root of 24 + the square root of 6? Simplify if possible.
Found 2 solutions by ewatrrr, Theo:
Answer by ewatrrr(24785)

 
Hi,Response previously posted

Answer by Theo(13342)

square root of (486) can be simplified as follows:
it is equivalent to:
square root of (486 * 1)
or:
square root of (54 * 9)
or:
square root of (6 * 9 * 9) which is equivalent to:
square root of (6 * 9^2)
since the square root of (x^2) is equal to x, this means that:
the square root of (6 * 9^2) is equal to 9 * square root of (6)
you can use your calculator to confirm that 9 * sqrt(6) is equivalent to sqrt(486)
the square root of (864) is not equivalent to the square root of (24) plus the square root of (6).
you can also use your calculator to confirm that statement is not true.
the arithmetic laws of square roots state:
sqrt(a*b) = sqrt(a) * sqrt(b)
sqrt(a/b) = sqrt(a) / sqrt(b)
those laws do not work with addition or subtraction.
sqrt(a+b) does NOT equal sqrt(a) + sqrt(b)
sqrt(a-b) does NOT equal sqrt(a) - sqrt(b)
here's some references that might help you to refresh your memory.
http://www.themathpage.com/alg/radicals.htm
http://www.themathpage.com/alg/simplify-radicals.htm
http://www.themathpage.com/alg/multiply-radicals.htm